Dave and Jane Wells have a new rectangle driveway. The perimeter of the driveway is 168 feet. The length is 12 feet longer then three times the width. What are the dimensions of the driveway?
Answer by nerdybill(7384)

Dave and Jane Wells have a new rectangle driveway. The perimeter of the driveway is 168 feet. The length is 12 feet longer then three times the width. What are the dimensions of the driveway?
.
Let w = width
then
3w+12 = length
.
2(w + 3w+12) = 168
2(4w+12) = 168
(4w+12) = 84
4w = 72
w = 13 feet (width)
.
Length:
3w+12 = 3(13)+12 = 39+12 = 51 feet